# Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions

### v2

@@ 6.3. Challenge 3: Safeguarding Co-adaptation
As advanced AI systems become increasingly complex, they present greater challenges for human interpretation and control. To address this, it is crucial to empower humans to detect and interpret AI misconduct on instrumental actions towards accomplishing its final goals. Additionally, mechanisms must be developed to enable human intervention and prevent power-seeking AI misconduct, thus maintaining a safe co-adaptation between AI and humans. Furthermore, advanced AI systems, even with sufficient interpretation capabilities, may intentionally mislead or disobey humans. It is possible for AI to stand by falsehoods, generate empty explanations for their answers, and produce outright fabrications that may appear plausible (Johnson and Iziev, 2022). Therefore, developing reliable interpretability mechanisms to validate the faithfulness and honesty of AI systems’ self-reported behaviors is also essential.
Examples of Potential Future Solutions. One potential research area aimed at empowering humans to interpret and control advanced AI behavior is the design of corrigible mechanisms that facilitate easy intervention and correction. This includes developing modular AI architectures where components can be independently adjusted or replaced, and creating robust override protocols that allow human operators to safely halt or redirect AI activities. These components should ideally be human-interpretable, enabling supervisors to conduct scenario testing when necessary. Another critical research direction lies in validating the faithfulness and honesty of AI interpretability. This could involve correlating AI behaviors with internal neuron activity signals, similar to how physiological indicators are measured in human polygraph tests (Association et al., 2004). Inspecting internal neuron indicators could support human assessment of the truthfulness of AI interpretations, thereby allowing them to accurately judge and prevent potential risky instrumental actions.
@@
Preventing future advanced AI systems from engaging in risky instrumental actions requires human capabilities to identify and mitigate such behaviors. Additionally, scalable solutions that can be widely implemented are essential for effectively supervising AI instrumental behavior across various applications and environments. Future research could potentially explore comprehensive training programs and the development of advanced diagnostic tools to achieve these goals.
@@
Examples of Potential Future Solutions. To enhance human capability in supervising AI instrumental actions, robust training and simulation environments are essential. These should offer scenario-based training with timely feedback loops, enabling supervisors to identify and manage various instrumental actions effectively. Additionally, creating interactive and intuitive dashboards is crucial for quickly recognizing and responding to potential risky instrumental goals. These dashboards should include effective data and model visualization, intervention techniques to highlight anomalies, and real-time alerts for prompt intervention. Furthermore, as deploying advanced AI at scale makes real-time oversight more challenging, developing advanced and autonomous monitoring tools is vital. These tools should learn the normal operating parameters of AI systems, flagging deviations that suggest risky instrumental actions immediately. By integrating training and simulation environments, interactive dashboards, and scalable diagnostic tools, humans can better manage AI instrumental risks, supporting better alignment with human values and control.
7. Implications and Limitations
In this section, we provide a summary of the key implications of our proposed framework, and discuss the challenges and limitations we encountered during the systematic literature review process.
7.1. Framework Implications and Applications
7.1.1. Cross-domain inspirations.
Achieving long-term human-AI alignment requires holistic, interdisciplinary efforts. This paper introduces a bidirectional human-AI alignment framework to create a shared vocabulary among diverse fields. Supported by over 400 multi-domain research papers, this framework enables researchers to integrate knowledge from other domains, enhancing their work and fostering synergetic collaboration. Particularly, AI researchers can use the human value taxonomy grounded in psychological theories (e.g., Section 5.2) to develop datasets and algorithms that incorporate a broader spectrum of human values. Additionally, by examining studies on individual and societal adaptations to AI alignment (e.g., Section 4.4.2), AI developers can create more practical and adaptive AI systems that evolve alongside human and societal changes. Conversely, HCI and social science researchers can leverage customizable AI alignment algorithms (e.g., Section 4.2) to design interactive systems that improve human interaction with AI. By using interaction techniques to specify human values (Section 4.1), these researchers can optimize interfaces or conduct surveys to extract human values, informing AI alignment development.
7.1.2. Foundation of Dynamic Alignment Evolution.
The proposed bidirectional human-AI alignment framework also serves a foundational role to support researchers to address the outlined future challenges and achieve long-term alignment. As elaborated in Section 6, within each direction, researchers can potentially address future challenges on the basis of existing alignment studies organized by the framework. For instance, to tackle the specification game, future AI researchers can use the comprehensive human value taxonomy outlined in the framework (Section 6) to develop methods for selecting the appropriate subset of these human values to be integrated into AI alignment. Additionally, we emphasize the importance of synergizing both directions to achieve long-term bidirectional human-AI alignment. To facilitate dynamic co-evolution (Section 6.2), we advocate for collaboration among AI, HCI, and social scientists to measure human and societal adaptations and incorporate these changes into the alignment of next-generation AI systems. This holistic approach can contribute to the evolution of AI in harmony with human values and societal needs.
7.2. Limitations
@@
One limitation of this work is the coverage of the sampled and filtered papers. The literature revolving around the topic of human-AI alignment is increasing rapidly, scattering in diverse venues across many domains. Therefore, instead of striving for an exhaustive collection of papers, this work focused on developing a holistic loop of bidirectional human-AI alignment framework by using essential research questions, dimensions and codes. Additionally, we are aware that our surveyed papers and the position of our team members primarily focus on computing-related domains, such as ML, NLP, and HCI. There are additional disciplines involved in alignment research, such as cognitive science, psychology, and STS (Science, Technology, and Society). However, our framework can naturally extend to these domains by incorporating them into the appropriate directions. Despite these limitations, we argue that this bidirectional human-AI alignment framework serves as foundational reference forfuture researchers examining existing literature that lies in the holistic loop of the alignment process, while preventing implicit assumptions or overlooked considerations, thereby facilitating a holistic understanding of the dimensions and factors that drive the developing of AI technology and human impacts. For future work, we plan to continue to refine our dimensions and codes to incorporate future literature relevant to the bidirectional human-AI alignment process.
8. Conclusion
@@
Acknowledgements.
We thank Eric Gilbert for his constructive feedback on human-centered insights on alignment, thank Eytan Adar for his valuable guidance on designing the interaction techniques for human-AI alignment, and thank Elizabeth F. Churchill for her insightful discussion on this manuscript. We also thank Michael S Bernstein, Denny Zhou, Cliff Lampe, and Nicole Ellison for their encouraging feedback on this work. We welcome researchers’ constructive discussions and interdisciplinary efforts to achieve long-term and dynamic human-AI alignment collaboratively in the future. This work was supported in part by the National Science Foundation under Grant No. IIS-2143529 and No. IIS-1949634.APPENDIX
Appendix A Author contributions
@@
A.1. Overall Author List and Contributions
@@
Appendix B Systematic Literature Review
B.1. Venues
HCI: CHI, CSCW, UIST, IUI;
NLP: ACL, EMNLP, NAACL, Findings
ML: ICLR, NeurIPS
Others: ArXiv, FAccT, AIES, and other related work
B.2. Keywords
Human: Human, User, Agent, Cognition, Crowd
Alignment: Align, Alignment
Value: Value, Principle
Trust: Trust, Trustworthy
Visualize: Visualization, Visualize
Explain: Interpretability, Explain, Understand, Transparent
Evaluation: Evaluate, Evaluation, Audit
Feedback: Feedback
Ethics: Bias, Fairness
B.3. Inclusion and Exclusion Criteria
RQ.1
RQ.2
RQ.3
RQ.4
RQ.5
RQ.6
Inclusion Criteria:
Exclusion Criteria:
Tasks art and design, emotion.
No human included
focus on English primarily focus on English as the main language;
Embedding-based explanation, evaluation, etc.
B.4. Collected Main Papers
HCI: (Xu and Zhang, 2023; Ding et al., 2023a; Srivastava et al., 2023; Wischnewski et al., 2023; Piorkowski et al., 2023; Lai et al., 2023; Weidele et al., 2023; Huang et al., 2023a; Mahmood et al., 2022; Guo et al., 2022; Mitchell et al., 2022; Zagalsky et al., 2021; Smith-Renner et al., 2020a, b; Yang et al., 2023a; Jo et al., 2023; Pinski et al., 2023; Cabitza et al., 2023; Hämäläinen et al., 2023; Ma et al., 2023a; Banovic et al., 2023; Vasconcelos et al., 2023; He et al., 2023a; Mok et al., 2023; Lee and Chew, 2023; Shen et al., 2023a; Schemmer et al., 2023; Kim et al., 2023b; He et al., 2023b; Wang et al., 2023f; Danry et al., 2023; Solyst et al., 2023a; Bennett et al., 2023; Lima et al., 2023a; Ashktorab et al., 2023; Wang and Yin, 2023; Jakesch et al., 2023; Solyst et al., 2023b; Jiang et al., 2023a; Prabhudesai et al., 2023; McNutt et al., 2023; Boggust et al., 2022; Tolmeijer et al., 2022; Zhang et al., 2022a; Chiang and Yin, 2022; Chen et al., 2022; Gajos and Mamykina, 2022; Lee and Rich, 2021; Liao and Sundar, 2021; Park et al., 2021a; Lima et al., 2021; Richardson et al., 2021; Van Berkel et al., 2021; Wang et al., 2021; Lee and Singh, 2021; Bandy, 2021; Rakova et al., 2021; Sovrano and Vitali, 2021; Szymanski et al., 2021; Eiband et al., 2021; Chromik et al., 2021; Wang and Yin, 2021; Hong et al., 2021; Gilad et al., 2021; Buçinca et al., 2021; Vereschak et al., 2021; Weisz et al., 2021; Liao et al., 2020; Lai et al., 2020; Ashktorab et al., 2020; Benedikt et al., 2020; Khadpe et al., 2020; Yin et al., 2019; Kocielnik et al., 2019; Holstein et al., 2019; Wang et al., 2019c; Cai et al., 2019; Wang et al., 2019b; Kapania et al., 2023; Lewicki et al., 2023; Varanasi and Goyal, 2023; Wong et al., 2023; Corvite et al., 2023; Kawakami et al., 2022; Sambasivan and Veeraraghavan, 2022; Boyd, 2021; Scheuerman et al., 2021; Ismail and Kumar, 2021; Madaio et al., 2020; Lima et al., 2020; McDonald and Pan, 2020; Zheng et al., 2023b; Sivaraman et al., 2023; Munyaka et al., 2023; Zhang et al., 2023c; Chen et al., 2023c; Kim et al., 2023d; He et al., 2023c; Lee et al., 2023c; Cabrera et al., 2023a; Lu et al., 2023b; Rogers and Crisan, 2023; Hou et al., 2023; Cabrera et al., 2023b; Petridis et al., 2023; Das et al., 2023; Wu et al., 2023d, e; Hemmer et al., 2023; Park et al., 2023; Cau et al., 2023; Zhang et al., 2023a; Wang et al., 2023e; Gebreegziabher et al., 2023; Liu et al., 2023b; Gao et al., 2023b; Cimolino and Graham, 2022; Zhang et al., 2022b; Wu et al., 2022a; Wang et al., 2022b; Suresh et al., 2022; Hong et al., 2022; Dang et al., 2022; Lai et al., 2022; Cunningham et al., 2021; Rietz and Maedche, 2021; Bansal et al., 2021; Jacobs et al., 2021; Jiang et al., 2021; Liu et al., 2021; Zhang et al., 2021a; Françoise et al., 2021; Cabrera et al., 2021; Wright et al., 2021; Desmond et al., 2021; Yang et al., 2020; Li, 2020; Choi et al., 2019; Wang et al., 2019a; Feng and Boyd-Graber, 2019; Amershi et al., 2019; Chung et al., 2022; Wei et al., 2023; Jörke et al., 2023; Long and Magerko, 2020; Petridis et al., 2024a; Chiang et al., 2023; Gu et al., 2024; Xu and Zhang, 2023)
NLP/AI
(Jiang et al., 2023b; Wu et al., 2023f; Pei et al., 2022; Dong et al., 2024; Bansal et al., 2023; Verma and Metcalf, 2023; Chen et al., 2023b; Zhou et al., 2023b; Isajanyan et al., 2024; Cao et al., 2024; Zhang et al., 2024; Ma et al., 2023b; Go et al., 2023; Binz and Schulz, 2023; Hu et al., 2023; Si et al., 2023a; Wang et al., 2023h; Guan et al., 2022; Suresh et al., 2023; Lahoti et al., 2023; Heinisch et al., 2023; Zhao et al., 2023b; Jeoung et al., 2023; Mckenna et al., 2023; Wu et al., 2023b; Nie et al., 2023; Zhou et al., 2023a; Hwang et al., 2023a; Dong et al., 2023a; Yan et al., 2023c; Zhan et al., 2023; Lin and Ng, 2023; Peng et al., 2023; Ye et al., 2024; Kang et al., 2023a; Li et al., 2023c; Liu and Zhu, 2023; Cai et al., 2023; Song et al., 2024; Dong et al., 2023b; Yu et al., 2023; Ethayarajh and Jurafsky, 2022; Gao et al., 2022; Feder et al., 2022; Jones and Steinhardt, 2022; Reddy et al., 2022; Bakker et al., 2022; Tseng et al., 2021; Patel and Pavlick, 2021; Prabhumoye et al., 2021; Park et al., 2021b; Paleja et al., 2021; Tian et al., 2020; Yan et al., 2023b; Chan et al., 2023; Lee et al., 2023b; Li et al., 2023b; Ding et al., 2023b; Schick et al., 2022; Cheng et al., 2022a; Zhang et al., 2021b; Vig et al., 2021; Tenney et al., 2020; Wallace et al., 2019; Goldfarb-Tarrant et al., 2019; Hosking et al., 2024; Lin et al., 2023b; Ruan et al., 2023; Felkner et al., 2023; Sharma et al., 2023; Ramezani and Xu, 2023; Sky et al., 2023; Ouyang et al., 2023; Goldfarb-Tarrant et al., 2023; Ma et al., 2023c; Rao et al., 2023; Huang et al., 2023c; Kiesel et al., 2022; Ammanabrolu et al., 2022; Sheng et al., 2021; Zhu et al., 2023; Klissarov et al., 2024; Zha et al., 2023; Feng et al., 2023; Joshi et al., 2023; Yao et al., 2023c; Kwon and Mihindukulasooriya, 2023b; Dhuliawala et al., 2023; Lee et al., 2023a; Tian et al., 2023a; Lin et al., 2023c; Qi et al., 2023; Kenny et al., 2022; Shen et al., 2022a; Khashabi et al., 2022; Voigt et al., 2022; Sun et al., 2022; Hase and Bansal, 2020; Gao et al., 2024; Liu et al., 2024b; Gou et al., 2024; Guo et al., 2024; Bradley et al., 2024; Li et al., 2024c, a; Ngo et al., 2024; Yuan et al., 2024; Jia et al., 2024; Xu et al., 2024; Tian et al., 2024; Petrak et al., 2024; Tian et al., 2023b; Xu et al., 2023a; Yang et al., 2023b; Wang et al., 2023a; Slobodkin et al., 2023; Yao et al., 2023b; Chung et al., 2023; Lu et al., 2023a; Oh et al., 2023; Li et al., 2023a; Kirk et al., 2023; Fleisig et al., 2023; Roit et al., 2023; Akyürek et al., 2023; Santy et al., 2023; Dai et al., 2024; Sicilia and Alikhani, 2023; Fei et al., 2023; Wang et al., 2023j; Fatemi et al., 2023; Selvam et al., 2023; Wang et al., 2023b; Xu et al., 2023b; Orlikowski et al., 2023; Gao et al., 2023a; Esiobu et al., 2023; Li et al., 2023d; Wang and Plank, 2023; Majumder et al., 2023; Shen et al., 2023b; Wang et al., 2023i; Chawla et al., 2023; Kim et al., 2023a, c; Kang et al., 2023b; Chen et al., 2023a; Fan et al., 2023; Shen et al., 2023c; Wang et al., 2023c; Si et al., 2023b; Wang et al., 2023g; Qian et al., 2023; Ouyang and Li, 2023; Loakman et al., 2023; Hu and Clune, 2023; Yuan et al., 2023; Hong et al., 2023; Sucholutsky and Griffiths, 2023; Zhang et al., 2023b; Khani and Ribeiro, 2023; Sarkar et al., 2023; Corvelo Benz and Rodriguez, 2023; Yao et al., 2023a; Yan et al., 2023a; Lin et al., 2023a; Torne Villasevil et al., 2023; Dubois et al., 2023; Sun et al., 2023; Wu et al., 2023c; Gupta et al., 2023; Mozannar et al., 2023; Lin et al., 2024; Hu et al., 2024b; Liu et al., 2023c; Mireshghallah et al., 2024; Dong and Ma, 2023; Dorner et al., 2023; Lu et al., 2022; Bonaldi et al., 2022; Geva et al., 2022; Wiegreffe et al., 2022; Li et al., 2022b; Khot et al., 2022; Dalvi et al., 2022; Ter Hoeve et al., 2022; Liu et al., 2022b; Nguyen et al., 2022; Ye and Durrett, 2022; Wu et al., 2022b; Liu et al., 2022a; Vodrahalli et al., 2022; Jin et al., 2022; Li et al., 2022a; Gerstgrasser et al., 2021; Ouyang et al., 2022; Guan et al., 2021; Arodi and Cheung, 2021; Boecking et al., 2020; Tay et al., 2020; Lertvittayakumjorn et al., 2020; Gao et al., 2020; Stiennon et al., 2020; Carroll et al., 2019; Sen et al., 2019; Zheng et al., 2023a; Zhao et al., 2023a; Lyu et al., 2023; Hwang et al., 2023b; Muthusamy et al., 2023; Lim et al., 2023; Hongru et al., 2023; Maghakian et al., 2023; Swazinna et al., 2023; Welch et al., 2022; Choi et al., 2023; Deng et al., 2023; Rafailov et al., 2023)
Others (Yao et al., 2023d; Sorensen et al., 2024; Wang et al., 2023l; Santurkar et al., 2023; Terry et al., 2023; Bai et al., 2022b; Ganguli et al., 2022; AlKhamissi et al., 2024; Bai et al., 2022a; Ziegler et al., 2019; Kim et al., 2024; Chen et al., 2024; Scheurer et al., 2023; Askell et al., 2021; Shilton, 2013; Huang et al., 2023b; Zhao et al., 2024; Wang et al., 2023k; Ashkinaze et al., 2024; Petridis et al., 2024b; Kaur et al., 2022; Zhou and Lee, 2024; Perry et al., 2023; Mozannar et al., 2024; Cheng et al., 2022b; Kazemitabaar et al., 2023; Hacker et al., 2023; Peschl et al., 2022; Shen et al., 2022b; Kwon and Mihindukulasooriya, 2023a; Sundar and Lee, 2022; Yildirim et al., 2022; Ma et al., 2024; Shaikh et al., 2024; Shen and Huang, 2020; Li et al., 2024b; Atreja et al., 2023; Lima et al., 2023b)

### v3

@@ 6.3. Safeguarding Co-adaptation
(i) As advanced AI systems become more complex, they present greater challenges for human interpretation and control. It is crucial to empower humans to detect and interpret AI misconduct and enable human intervention to prevent power-seeking AI behavior. Research should focus on designing corrigible mechanisms for easy intervention and correction, including modular AI architectures and robust override protocols that allow human operators to halt or redirect AI activities. These components should be human-interpretable, enabling scenario testing. (ii) Furthermore, advanced AI systems may intentionally mislead or disobey humans, generating plausible fabrications (Johnson and Iziev, 2022). Developing reliable interpretability mechanisms to validate the faithfulness and honesty of AI behaviors is essential. This includes correlating AI behaviors with internal neuron activity signals, akin to physiological indicators in human polygraph tests (Association et al., 2004). Inspecting these indicators can help humans assess the truthfulness of AI interpretations and prevent risky actions.
@@
(i) Preventing advanced AI from engaging in risky actions requires robust human supervision. Essential steps include developing training and simulation environments with scenario-based exercises and timely feedback, and creating interactive dashboards for real-time monitoring. These dashboards should feature effective data visualization, anomaly detection, and prompt alert systems for immediate intervention. (ii) Scalable solutions are needed for supervising AI across various applications. Real-time oversight becomes more challenging with widespread AI deployment, necessitating advanced autonomous monitoring tools. These tools should learn normal AI behavior and flag deviations immediately. Integrating training environments, interactive dashboards, and scalable diagnostic tools will enhance human ability to manage AI risks, ensuring better alignment with human values.
7. Limitations and Conclusion
@@
One limitation of this work is the scope of the sampled and filtered papers. The rapidly growing literature on human-AI alignment spans diverse venues across many domains. Instead of an exhaustive collection, we focused on developing a holistic bidirectional human-AI alignment framework using essential research questions, dimensions, and codes. Our surveyed papers and team members primarily focus on computing-related fields like ML, NLP, and HCI, though alignment research also involves disciplines like cognitive science, psychology, and STS (Science, Technology, and Society). Our framework can naturally extend to these areas as needed. Despite these limitations, we believe our bidirectional human-AI alignment framework serves as a foundational reference for future researchers.
@@
APPENDIX
Appendix A Systematic Literature Review
A.1. Systematic Literature Review Process
A.1.1. Identification and Screening with Keywords.
We started with papers published in the AI-related domain venues (including NLP, HCI, and ML fields) beginning from the advent of general-purpose generative AI to present, i.e., primarily between January, 2019 and January, 2024 (see details in Appendix A.2). We retrieved 34,213 papers in the initial Identification stage. Further, we collectively defined a list of keywords (see details in Appendix A.3) and screened for papers that included at least one of these keywords (e.g., human, alignment) or their variations either in the title or abstract. We included 2,136 papers in this Screening stage.
A.1.2. Assessing Eligibility with Criteria
We further filtered the 2,136 papers based on explicit inclusion and exclusion criteria, i.e., the Eligibility stage. Our criteria revolved around six research questions that we collectively identified to be most pertinent to the topic, including 1) what essential human values have been aligned by some AI models? 2) how did we effectively quantify or model human values to guide AI development? 3) what strategies have been employed to integrate human values into the AI development process? 4) how did existing studies improve human understanding and evaluation of AI alignment? 5) what are the practices for designing interfaces and interactions that facilitate human-AI collaboration? 6) How have AI been adapted to meet the needs of various human value groups? We included papers that could potentially answer any of these questions. Further, based on the scope in Section 3.4, we excluded papers that did not meet our inclusion criteria. This resulted in a final corpus of 411 papers, which were analyzed in detail using qualitative coding (see Appendix A.4 for more details).
A.1.3. Qualitative Code Development.
Referring to the code development process in Lee et al. (2024), we first conducted qualitative coding for each paper by identifying relevant sentences that could answer the above research questions, and entering short codes to describe them into a codebook. We iteratively coded relevant sentences from each paper through a mix of inductive and deductive approaches, which allowed flexibility to expand, modify or change the driving research questions based on our learnings as we went through the process. To ensure rigor in our coding process, two authors coded each paper. The first author independently annotated all papers after reviewing the paper abstracts and introductions. Twelve team members each annotated a subset of the paper corpus. Our corpus includes papers from different domains (e.g., HCI, NLP and ML). Therefore, we divided the authors into HCI and NLP/ML222Note that NLP and ML are two different domains, we combine them together for the purposes of literature review analysis since they both work on developing and evaluating AI technologies. teams and assigned the papers accordingly based on expertise. All team members coded each of their assigned papers to answer all six questions (if applicable) introduced above.
A.1.4. Framework Development and Rigorous Coding.
After developing annotations, all authors collaborated to create the bidirectional human-AI alignment framework by integrating the annotations within each of the codes. The initial version of the framework was proposed by the author who reviewed all papers. This framework furthermore underwent iterative improvement through: 1) discussions with all team members involved in paper coding, and 2) revisions based on feedback from the project advisors. Additionally, we strengthened the framework by reviewing papers from the AI Ethics conferences (including FAccT and AIES), and related work of the collected papers that covered other domains such as psychology and social science. We further added missing codes and papers to ensure comprehensive coverage (see Appendix A.2 for details). The final bidirectional human-AI alignment framework, with detailed topologies, is presented in Section 4. Following the framework’s finalization, we conducted another separate coding process to annotate whether each paper investigated dimensions within our framework. Two authors independently coded each paper.333The joint probability of agreement for the paper annotations was 0.78. These codes were then used to perform quantitative and qualitative analyses, as presented in Section 5.
A.2. Venues
HCI: CHI, CSCW, UIST, IUI;
NLP: ACL, EMNLP, NAACL, Findings
ML: ICLR, NeurIPS
Others: ArXiv, FAccT, AIES, and other related work
A.3. Keywords
Human: Human, User, Agent, Cognition, Crowd
Alignment: Align, Alignment
Value: Value, Principle
Trust: Trust, Trustworthy
Visualize: Visualization, Visualize
Explain: Interpretability, Explain, Understand, Transparent
Evaluation: Evaluate, Evaluation, Audit
Feedback: Feedback
Ethics: Bias, Fairness
A.4. Inclusion and Exclusion Criteria
RQ.1
RQ.2
RQ.3
RQ.4
RQ.5
RQ.6
Inclusion Criteria:
Exclusion Criteria:
Tasks art and design, emotion.
No human included
focus on English primarily focus on English as the main language;
Embedding-based explanation, evaluation, etc.
Appendix B Author contributions
@@
B.1. Overall Author List and Contributions